to pay for an $18,600 car elsa made a down payment of $4500 and tok oout a loan for the rest. On the loan she paid monthly payments of $312.10 for 4 years What was the total amount Elsa ended up paying for the car (including the down payment and monthly payment? How much interest did Elsa pay on the loan


1
Expert's answer
2021-09-02T00:31:29-0400

Amount paid as monthly payments "=(12\u00d74\u00d7312.10)=14,980.8" dollars


Total amount Elsa paid for the car "=(14980.8+4500)=19,480.8" dollars


Interest Elsa paid on loan "=(19480.8-18600)=880.8" dollars